This is a ic Entry question/It is worth 3 points/You have u 22 Question (3 points) A total of 90.00 kg of a fertilizer that contains 30.00 % nitrogen. 2nd attempt Part 1 (1 point) Feedback How much nitrogen is in the fertilizer? kg N 22:30

media%2Fac9%2Fac98e335-1e0c-41bf-99b9-4a

I'm lost on this question....
Answer #1

Part 1:

Total weight of fertilizer = 90 Kg

Nitrogen contain = 30%

So, 30% of 90 kg = (30/100) x 90 Kg
                              = 27 Kg

So, it contains 27 Kg of Nitrogen.

Part 2:

Total weight of fertilizer = 90 Kg

15% is washed into the river.

So, 15% of 90 kg = (15/100) x 90 Kg
                              = 13.5 Kg

Now, Nitrogen contained is 30%

So, 30% of 13.5 kg = (30/100) x 13.5 Kg
                                = 4.05 Kg

Part 3:

1 Kg = 1000 g

1 g = 1000 mg

So, 1 Kg = 106 mg

Or 4.05 Kg = 4.05 x 106 mg
                   = 4050000 mg